Andretti Company has a single product called a Dak. The company normally produces and sells 82,000 Daks each year at a sellinReg 1A Reg 1B Reg 2 Reg 3 Req 4A to 4C Req 4D Reg 5 Assume that Andretti Company has sufficient capacity to produce 106,600 DReq 1A Reg 1B Req 2 Reg 3 Req 4A to 4C Req 4D Reg 5 Assume again that Andretti Company has sufficient capacity to produce 106Req 1A Reg 1B Reg 2 Req3 Req 4A to 4C Req 4D Reg 5 The company has 900 Daks on hand that have some irregularities and are theReq 1A Reg 1B Reg 2 Reg 3 Reg 4A to 4C Req 4D Reg 5 Due to a strike in its suppliers plant, Andretti Company is unable to puReq 1A Reg 1B Reg 2 Reg 3 Req 4A to 4C Reg 4D Req 5 An outside manufacturer has offered to produce 82,000 Daks and ship them
Answer #1

1). Financial advantage (disadvantage) of investing an additional $120,000 in fixed selling expenses.

Present capacity = 106,600 daks

Currently making = 82,000 daks

Additional capacity can be used = 82000 * 30% = 24,600 daks

Incremental contribution from additional sales :

Selling price = $60

Less: Variable costs:

Direct materials =$8.50

Direct labor = $10.00

Variable manufacturing overhead = $2.70

Variable selling expenses = $1.70

Incremental contribution per unit = $37.10

Total Incremental contribution for 24,600 daks = 24,600 * $37.10 = $912,660

Less: Incremental Fixed selling expenses = $130,000

Financial advantage (disadvantage) = $782,660

1b). Yes, additional investment is justified because there is a financial advantage of $782,660.

2). Calculation of Breakeven price per unit:

Total cost per unit for this outside order:-

Direct materials =$8.50

Direct labor = $10.00

Variable manufacturing overhead = $2.70

Variable selling expenses = $1.60

Import duties = $2.70

Permits and licences ($19,680 / 24600) = $0.8

Total cost per unit = $26.30

Hence the breakeven price per unit is $26.30

3). As the 900 daks are already been made and cannot be sold due to irregularities on normal price, its cost of production is sunk cost now because cost are already been incurred and cannot be recovered. The only unit cost which is relevant to these units is variable selling cost i.e. $1.70 per unit.

4).

a). Total contribution margin will Andretti forgo if it closes the plant for two months.

Contribution margin per unit = $37.10

Total contribution margin it will forgo = (82000 *25% * 2/12 )* $37.10 = $126,758.33

b). Total fixed cost will the company avoid if it closes the plant for two months.

Fixed manufacturing overhead ($574,000 * 65% * 2 /12) = $62,183.33

Fixed selling expenses ($246,000 * 20% * 2/12) = $8,200

Total avoidable fixed cost = $70,383.33

c). Financial advantage (disadvantage) of closing the plant for the two-month period:

Avoidable fixed cost = $70,383.33

Less: Loss of contribution margin = $126,758.33

Financial advantage (disadvantage) = $ (56,375)

d). Andretti should not close the plant for 2 months because there is financial disadvantage.

5). Andretti's avoidable cost per unit that it should compare to the price quoted by the outside manufacturer.

Direct materials =$8.50

Direct labor = $10.00

Variable manufacturing overhead = $2.70

Fixed manufacturing overhead ($7.0 * 30%) = $2.10

Variable selling expenses ($1.70 * 1/3) = $0.57

Total avoidable cost per unit = $23.87